Systems Staff Meeting Notes 11/09/98

Systems Staff Meeting Notes
November 9, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng

Announcements
  1. Please get your Homecoming Picnic ticket request forms to Debra.
  2. Ashley Jones should officially begin his position as Network Coordinator on Monday, November 23.
  3. The title change for Lee's old position has finally gone through - should be posted soon & filled (hopefully) by the end of the year.
  4. Winston's Programmer Analyst position - an ad will run in this Sunday's paper.
  5. Preservation Theft - Police think the theft of 3 CPU's over the weekend was extremely routine. They did, at least, dust for fingerprints. They left the monitors, but took 3 CPU's from the loft area of Preservation. There is always the chance that the thieves will return, so beware.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Disk upgrades - received 2 Gigs instead of 6.8 Gigs.
  2. MSL - Bobbie notes that they have 8-10 Mac Performas stored there. Can we just have PK Yonge pick the machines up from there? Yes, just need to physically go over there to work it out and keep the records straight - make sure they're not swapped out by MSL at the last minute.
Liaisons
  1. New Liaison Orientation is scheduled for November 18, 19 & 20. The November 19th session will be held at the regular Liaison time, 2:00-3:00 PM and is mandatory for the new liaisons only. In addition, they will need to attend one of the other meetings on either November 18th or 20th. Those meetings will be held from 1:00-2:00 PM. All classes will be conducted in room 148 - the electronic classroom in Library West. Bill is trying to streamline the paperwork for this session.
  2. Liaisons are in the middle of performing the Outlook changeovers. This process doesn't always work as it should. Debra will post the instructions on the Liaison page of our website.
Networking
  1. The change to Smathers NT2 should occur by the end of the month. After reading the documentation, we've decided to bring the new server up, update it to 5.5, then integrate the files. Smathers NT7 will be the second mail server - both will run simultaneously so that mail will not have to go down. All clients point to NT2 - they will have to be changed to NT7. File Shares on NT2 will remain. New server will be called NT7, this will make the back off easier, won't have to deal with tapes.
  2. ATM failure at CORE - 1:00 AM this past Saturday. CSE catalyst failure - IBM 8265's at the middle are cause.
Software Projects Status
  1. Mellon - Winston working on free text searches on Abstracts in the database - evaluating third part product.
  2. Personnel - Winston delivered the new interface to Erika, she is currently evaluating it.
  3. OPS Tracking - software has been installed on Resource Services machines, some staff have been trained on its use, but no data has been input yet.
  4. Timer software - John Ashcraft needed some changes made, but he is now satisfied. Project is finished.
  5. Newspaper Project - Laverne & Susan need to define the fields & structure - will meet with Brian Burton.
  6. Barbara Mattson is working on LASSIE with Brian.
DP Coordinator
  1. Sharon Wheeler has not gotten back with Debra regarding the Compustat tapes.
  2. Four SPOT tapes received last Friday afternoon - will send originals to NERDC to be copied.
Calendar
  1. Wednesday, November 11 - Veteran's Day.
  2. Brian Burton out November 20, 24 & 25.
  3. Thanksgiving Holiday - November 26 & 27.


Back to the Systems Homepage

Last updated February 24, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu